How do you work this problem: 10-5x=1x+18
Answer by laurien(6)

Hello,
You need to bring variables to one side first so 10-5x = 1x + 18
+ 5x=+5x
______________
10 = 6x + 18
now you can bring the constants to the other side -18
-18 = 6x
___________________
10-18 = -8 -8 = 6x/6
now just divide by the 6 /6 = x
so x =-8/6 = -4/3 = -1 1/3..however you need to write your solution
Think Math........:)